### v1

@@ 4.1 Data Sets
We adopt the medium and hard protein optimization benchmarks on AAV and GFP from Kirjner et al. (2023). Given the full data set $\mathcal{S}^{*}$, the task difficulty is determined by (i) the fitness percentile range of the considered sequences ($20-40$ for medium, $<30$ for hard) and (ii), the required gap of mutations to reach any sequence of the 99th fitness percentile of $\mathcal{S}^{*}$ (a gap of $6$ mutations for medium, $7$ mutations for hard), see Table 6. Essentially, this gives four different tasks described in Table 1, that each only sees a limited number of $N$ sequences within a limited fitness range. The idea of protein fitness optimization is to enhance these sequences to higher, previously unseen fitness values.
For in-silico evaluation of the generated sequences with the oracle $g_{\psi}$, we use the median normalized fitness, diversity and novelty following Jain et al. (2022), see Section A.2. While diversity and novelty are reported, no definitive higher or lower values are considered superior for a sequence. Note that $y_{\mathrm{min}}$ and $y_{\mathrm{max}}$ from the entire data set $\mathcal{S}^{*}$ are used for both GFP and AAV to normalize fitness scores to $[0,1]$.
@@ 4.2 Implementation Details
For training, we start by learning the VAE to compress the sequence token input space ($d=28$ and $d=237$ for AAV and GFP) to $l=16$ and $l=32$. A learning rate of $0.001$ with a convolutional architecture and $\beta\in\{0.01,0.001\}$ for AAV and GFP is used for training the encoder $\mathcal{E}$ and decoder $\mathcal{D}$ in Equation 4, also see Section B.2. To learn the prior distribution of embedded latent sequences $z=\mathcal{E}(x)$ using flow matching, the 1D CNN commonly used for denoising diffusion probabilistic models (DDPMs)  [^1] https://github.com/lucidrains/denoising-diffusion-pytorch is employed. A learning rate of $-55e\text{-}$ 5 and a batch size of $1024$ were used to train $v_{\theta,t}$ for $1000$ epochs. The predictors $g_{\phi}$ and $g_{\tilde{\phi}}$ are re-used without any further training Kirjner et al. (2023).
@@
At inference time, we follow the procedure outlined in Algorithm 1 to generate samples. We start from $z_{0}\sim\mathcal{N}(0,I)$ and use $K=32$ ODE steps to integrate the learned flow until we obtain $z_{1}$. To optimize sequence fitness, the condition $y=1$ is selected for all samples. The parameters $\alpha_{t}$ and $J$ are determined via a hyperparameter search, as detailed in Section B.4. After generating 512 samples $z_{1}$ to encourage sampling from the entire learned distribution, they are decoded using $x=\mathcal{D}(z_{1})$. Potential duplicates are then filtered out, and the top-k ($k=128$) samples, ranked by the predictor ($g_{\phi}$ or $g_{\tilde{\phi}}$, respectively), are selected.
@@
We use $g_{\phi}$ and compare VLGPO to GWG (which also uses the same trained predictor), and the identical smoothed predictor $g_{\tilde{\phi}}$ in VLGPO to GGS Kirjner et al. (2023). We benchmark against the respective baselines, namely GFlowNets (GFN-AL) Jain et al. (2022), model-based adaptive sampling (CbAS) Brookes et al. (2019), greedy search (AdaLead) Sinai et al. (2020), Bayesian optimization (BO-qei) Wilson et al. (2017), conservative model-based optimization (CoMS) Trabucco et al. (2021) and proximal exploration (PEX) Ren et al. (2022). Moreover, we investigate the performance of the recently introduced gradient-guided walk-jump sampling algorithm (gg-dWJS), which extends the walk-jump sampling framework originally developed for antibody sequence design Ikram et al. (2024); Frey et al. (2023). Because the available source code was not directly compatible with the data sets in Table 1, we adapted our existing model architectures for implementation. We then performed a grid search over sampling parameters, followed by top-k sampling for each task, to ensure a fair comparison.
@@ 4.3 Results
When it comes to diversity and novelty, there is no clear definition of what is better – higher or lower scores Kirjner et al. (2023). In general, we observe that reported values of VLGPO are in line with competing methods. Interestingly, the sequences in the 99th fitness percentile, which are on the top end of high-fitness sequences, show a diversity of 4.73 for GFP and 5.23 for AAV, which are approximately comparable to our results in Table 2 and Table 3, except for AAV hard. In practice, the smoothed predictor $g_{\tilde{\phi}}$ reduces the diversity within the generated sequences for both GGS and VLGPO. Additionally, more samples tend to collapse to the same decoded sequence, likely due to the smoother, less distinct gradients.
@@ 4.4 Fitness Extrapolation
Table 1 illustrates the limited fitness range of the sequences in each task, whereas protein fitness optimization aims to sample sequences with higher fitness values. Because the available data set is small, extrapolating to these high-fitness sequences during sampling becomes especially challenging. The following experiment examines how much the oracle-evaluated fitness $y_{\mathrm{gt}}$ (provided by $g_{\psi}$) deviates from the target fitness $y$, which is used as an input to the sampling in Algorithm 1. Thereby, we compare two approaches: our variational method VLGPO and a directly learned posterior $\mathrm{p}(x|y)$ via fitness-conditioned flow matching $v_{\theta,t}(z_{t},y)$. The learned posterior does not need the predictor $g_{\phi}$ for classifier guidance, as it learns the conditional distribution in the latent space end-to-end. For a fair comparison, the raw output without top-k sampling is applied. The results are depicted in LABEL:fig:variational_vs_posterior for GFP and AAV hard (see Section B.1 for AAV medium).
Due to the sparse data availability and the limited fitness range, the evaluated fitness $y_{\mathrm{gt}}$ cannot be expected to precisely follow the required fitness $y$. Nevertheless, the results in LABEL:fig:variational_vs_posterior clearly demonstrate the advantage of classifier guidance in VLGPO, whereas the direct posterior fails to effectively exploit the given fitness condition. This gap is most evident in the GFP hard task, where the training data’s fitness range is only in $[0.0,0.1]$, making it difficult for the learned posterior to extrapolate to higher fitness values. In contrast, VLGPO leverages the additional gradient information from the predictor $g_{\phi}$, thereby overcoming these limitations in higher-fitness regions.
@@ 4.5 Ablation Studies
We conduct an ablation study on the influence of manifold constrained gradients in sampling (Line 7, Algorithm 1). The results in Table 4 and Table 5 demonstrate the improvement gained by estimating $\hat{x}_{1}$ to compute the gradient of the likelihood term. Further, the performance of the directly learned posterior $\mathrm{p}(x|y)$ was investigated.While this approach still performs well for the medium tasks, it shows a larger performance drop on GFP (hard), indicating that it cannot match the explicit likelihood guidance provided by the fitness predictor in our variational method VLGPO.
@@ 5 Discussion
A limiting factor of our method lies in its hyperparameter tuning requirements. We observe that hyperparameter selection becomes more critical for challenging tasks such as GFP (hard), while it remains stable and robust for other tasks. Another important point is our reliance on in-silico evaluation, where we follow Kirjner et al. (2023) in using a trained oracle as the ground truth. This oracle was trained on a significantly larger data set than the tasks presented in Table 1 and can therefore be expected to serve as a good estimator. Nevertheless, experimental validation could provide valuable insights into the applicability of VLGPO.
@@ Appendix A Additional Methods
A.1 Data set
Table 6 describes the criteria that define the medium and hard tasks established by Kirjner et al. (2023), in terms of fitness percentile range and the required gap of mutations from the 99th fitness percentile of the entire data set $\mathcal{S}^{*}$.
Table 6: Task definition.
Task Range % Gap
Medium 20-40 6
Hard $<$ 30 7
@@ A.2 Metrics
For the evaluation of the generated sequences, we use median fitness, diversity and novelty as described in  Kirjner et al. (2023); Jain et al. (2022). The sequences are assessed by the oracle $g_{\psi}$ that was trained on the full data set $\mathcal{S}^{*}$ with minimum and maximum fitness values $y_{\mathrm{min}}$ and $y_{\mathrm{max}}$. The median normalized fitness of sampled sequences $x\in\mathcal{X}$ is computed using
@@ B.2 Variational Autoencoder
We train a VAE for each task to achieve at least $80\%$ reconstruction accuracy on a validation subset, balancing the latent prior with $\beta$. Table 7 shows these results. The lower accuracy in AAV compared to GFP is due to larger effects of mutations on shorter proteins (sequence length $d$). Since the data sets (see Table 1) are very small, the Kullback-Leibler (KL) divergence in Equation 4 is crucial to ensure a Gaussian latent distribution. We can also sample from the VAE and evaluate predicted fitness via $g_{\psi}$ (Table 7), although median normalized fitness is lower than in Table 1. This is expected, given the sparse latent space and the “hole” problem of VAEs Rezende & Viola (2018).
@@ B.4 Sampling Parameters
The determination of the hyperparameters $\alpha_{t}$ and $J$ in VLGPO sampling is performed through a grid search across different tasks. We keep $\alpha_{t}$ constant for all steps $t$. The resulting heatmaps for these two parameters are shown in LABEL:fig:sampling_sup, where for all tasks, band-like structures are evident, which are either unimodal or bimodal in shape. Overall, there appears to be a broad range from which suitable parameter choices for $\alpha_{t}$ and $J$ can be made for GFP (medium) and AAV (hard) and (medium). On the other hand, GFP (hard) displays distinctly different characteristics, with a wider fitness value range, suggesting that hyperparameter selection plays a more critical role when tackling a more difficult task. Due to considerably limited diversity of generated sequences at the maxima and their proximity for $\alpha_{t}$ and $J$ with GFP (hard), we choose significantly lower values $\alpha_{t}=0.02$ and $J=5$ to avoid compromising diversity, as fitness drops substantially when these hyperparameters grow too large. For the other tasks, we use the hyperparameter settings obtained from the grid search experiments, i.e., $\alpha_{t}\in\{0.97,1.2,0.56\}$ and $J\in\{39,19,37\}$ for AAV (medium), AAV (hard) and GFP (medium). Nevertheless, we do not observe major differences if these hyperparameters are adjusted slightly.
@@
Figure 9: Median fitness (left) and diversity (right) for all four tasks depending on employed ODE steps $K$ in sampling using Algorithm 1.

### v2

@@ 4.1 Data Sets
We adopt the medium and hard protein optimization benchmarks on AAV and GFP from (Kirjner et al., 2023). Given the full data set $\mathcal{S}^{*}$, the task difficulty is determined by (i) the fitness percentile range of the considered sequences ($20-40$ for medium, $<30$ for hard) and (ii), the required gap of mutations to reach any sequence of the 99th fitness percentile of $\mathcal{S}^{*}$ (a gap of $6$ mutations for medium, $7$ mutations for hard), see Table 1.
Table 1: Task definition.
Task Range % Gap
Medium 20-40 6
Hard $<$ 30 7
Together, this results in four different tasks described in Table 2, each of which only sees a limited number of $N$ sequences in a limited fitness range. The idea of protein fitness optimization is to enhance these sequences to higher, previously unseen fitness values. The setting reflects realistic scenarios in terms of data set sizes. The diversity (Section A.2) in the training data sets of the four tasks is quite high: for GFP (medium) it is 14.5, for GFP (hard) 16.3, for AAV (medium) 15.9, and for AAV (hard) 18.4. On the other hand, the diversity within the top-performing (99th percentile) sequences of the full data set $\mathcal{S}^{*}$ is 4.73 for GFP and 5.23 for AAV.
@@
For in-silico evaluation of the generated sequences with the oracle $g_{\psi}$, we use the median normalized fitness, diversity and novelty following (Jain et al., 2022), see Section A.2. The oracle was trained on the complete DMS data with 56,086 mutants for GFP and 44,156 mutants for AAV. While diversity and novelty are reported in evaluation, no definitive higher or lower values are considered superior for a sequence. Note that $y_{\mathrm{min}}$ and $y_{\mathrm{max}}$ from the entire data set $\mathcal{S}^{*}$ are used for both GFP and AAV to normalize fitness scores to $[0,1]$.
@@ 4.2 Implementation Details
For training, we start by learning the VAE to compress the sequence token input space ($d=28$ and $d=237$ for AAV and GFP) to $l=16$ and $l=32$. A learning rate of $0.001$ with a convolutional architecture and $\beta\in\{0.01,0.001\}$ for AAV and GFP is used for training the encoder $\mathcal{E}$ and decoder $\mathcal{D}$ in Equation 4, also see Section B.2. To learn the prior distribution of embedded latent sequences $z=\mathcal{E}(x)$ using flow matching, the 1D CNN commonly used for denoising diffusion probabilistic models (DDPMs)  [^2] https://github.com/lucidrains/denoising-diffusion-pytorch is employed. A learning rate of $-55e\text{-}$ 5 and a batch size of $1024$ were used to train $v_{\theta,t}$ for $1000$ epochs. All VAE and flow matching models are always trained under the limited‐data setting as listed in Table 2.
@@
At inference time, we follow the procedure outlined in Algorithm 1 to generate samples. The predictors $g_{\phi}$ and $g_{\tilde{\phi}}$ (for details, see Section A.1) are re-used without any further training (Kirjner et al., 2023). We start from $z_{0}\sim\mathcal{N}(0,I)$ and use $K=32$ ODE steps to integrate the learned flow until we obtain $z_{1}$. To optimize sequence fitness, the condition $y=1$ is selected for all samples. The parameters $\alpha_{t}$ and $J$ are determined via a hyperparameter search, as shown in LABEL:fig:sampling and discussed in detail in Section B.4. Ultimately, they modulate the trade-off between increasing fitness and maintaining diversity. After generating 512 samples $z_{1}$ to encourage sampling from the entire learned distribution, they are decoded using $x=\mathcal{D}(z_{1})$. Potential duplicates are then filtered out, and the top-k ($k=128$) samples, ranked by the predictor ($g_{\phi}$ or $g_{\tilde{\phi}}$, respectively), are selected. Note that also the predictors $g_{\phi}$ and $g_{\tilde{\phi}}$ for each setting, used for classifier guidance and for ranking the samples, are trained only on the data sets listed in Table 2.
@@
For evaluation of the generated samples, we use the oracle $g_{\psi}$ as in-silico fitness estimate, see Section A.1. The oracle is directly sourced from (Kirjner et al., 2023) and is the only model that was trained using the entire data $\mathcal{S}^{*}$. For sampling following Algorithm 1, we use $g_{\phi}$ and compare VLGPO to GWG (which also uses the same trained predictor), and the identical smoothed predictor $g_{\tilde{\phi}}$ in VLGPO to GGS (Kirjner et al., 2023). We benchmark against the respective baselines, namely GFlowNets (GFN-AL) (Jain et al., 2022), model-based adaptive sampling (CbAS) (Brookes et al., 2019), greedy search (AdaLead) (Sinai et al., 2020), Bayesian optimization (BO-qei) (Wilson et al., 2017), conservative model-based optimization (CoMS) (Trabucco et al., 2021) and proximal exploration (PEX) (Ren et al., 2022). Moreover, we investigate the performance of the recently introduced gradient-guided walk-jump sampling algorithm (gg-dWJS), which extends the walk-jump sampling framework originally developed for antibody sequence design (Ikram et al., 2024; Frey et al., 2023). Because the available source code was not directly compatible with the data sets in Table 2, we adapted our existing model architectures for implementation. We then performed a grid search over sampling parameters, followed by top-k sampling for each task, to ensure a fair comparison.
@@ 4.3 Results
When it comes to diversity and novelty, there is no clear definition of what is better – higher or lower scores (Kirjner et al., 2023). In general, we observe that reported values of VLGPO are in line with competing methods. Moreover, the variations between different methods in terms of diversity and novelty are significantly larger in GFP than in AAV. That discrepancy may arise due to GFP sequences being longer, thus representing a sparser and higher-dimensional search space that makes the protein harder to optimize. Interestingly, the sequences in the 99th fitness percentile, which are on the top end of high-fitness sequences, show a diversity of 4.73 for GFP and 5.23 for AAV, which are approximately comparable to our results in Table 3 and Table 4, except for AAV hard. In practice, the smoothed predictor $g_{\tilde{\phi}}$ reduces the diversity within the generated sequences for both GGS and VLGPO. Additionally, more samples tend to collapse to the same decoded sequence, likely due to the smoother, less distinct gradients.
@@ 4.4 Fitness Extrapolation
Table 2 illustrates the limited fitness range of the sequences in each task, whereas protein fitness optimization aims to sample sequences with higher fitness values. Because the available data set is small, extrapolating to these high-fitness sequences during sampling becomes especially challenging. The following experiment examines how much the oracle-evaluated fitness $y_{\mathrm{gt}}$ (provided by $g_{\psi}$) deviates from the target fitness $y$, which is used as an input to the sampling in Algorithm 1. Thereby, we compare two approaches: our variational method VLGPO and a directly learned posterior $\mathrm{p}(x|y)$ via fitness-conditioned flow matching $v_{\theta,t}(z_{t},y)$. The learned posterior does not need the predictor $g_{\phi}$ for classifier guidance, as it learns the conditional distribution in the latent space end-to-end. For a fair comparison, the raw output without top-k sampling is applied. The results are depicted in LABEL:fig:variational_vs_posterior for GFP and AAV hard.
Due to the sparse data availability and the limited fitness range, the evaluated fitness $y_{\mathrm{gt}}$ cannot be expected to precisely follow the required fitness $y$. Nevertheless, the results in LABEL:fig:variational_vs_posterior clearly demonstrate the advantage of classifier guidance in VLGPO, whereas the direct posterior fails to effectively exploit the given fitness condition. This gap is most evident in the GFP hard task, where the training data’s fitness range is only in $[0.0,0.1]$, making it difficult for the learned posterior to extrapolate to higher fitness values. In contrast, VLGPO leverages the additional gradient information from the predictor $g_{\phi}$, thereby overcoming these limitations in higher-fitness regions. This highlights the advantages of using a separate classifier for guidance in domains where classifier-free guidance struggles to produce high-fitness sequences.
@@ 4.5 Ablation Studies
We conduct an ablation study on the influence of manifold constrained gradients in sampling (Line 7, Algorithm 1). The results in Table 5 and Table 6 demonstrate the improvement gained by estimating $\hat{x}_{1}$ to compute the gradient of the likelihood term.
Further, the performance of the directly learned posterior $\mathrm{p}(x|y)$ was investigated. While this approach still performs well for the medium tasks, it shows a larger performance drop on GFP (hard), indicating that it cannot match the explicit likelihood guidance provided by the fitness predictor in our variational method VLGPO.
@@ 5 Discussion
A limiting factor of our method lies in its hyperparameter tuning requirements. We observe that hyperparameter selection becomes more critical for challenging tasks such as GFP (hard), while it remains stable and robust for other tasks. Additionally, the restriction of the benchmark to only AAV and GFP is a limitation that should be addressed in future work. Conceptually, VLGPO as well as competing approaches can be extended to other proteins from FLIP (Dallago et al., 2021) or ProteinGym (Notin et al., 2023).
Another important point is our reliance on in-silico evaluation, where we follow (Kirjner et al., 2023) in using a trained oracle as the ground truth. This oracle was trained on a significantly larger data set than the tasks presented in Table 2 and can therefore be expected to serve as a good estimator. Nevertheless, actual experimental validation could provide valuable insights into the applicability of VLGPO. A complementary direction may be to design more idealised, synthetic in-silico benchmarks that are nevertheless good proxies for protein design (Stanton et al., 2024). Moreover, additional metrics such as folding confidence or structural stability could be added to obtain a more complete perspective beyond the fitness score (Johnson et al., 2023).
Acknowledgements
This work was supported by the University Research Priority Program (URPP) Human Reproduction Reloaded of the University of Zurich.
Impact Statement
This paper presents a computational method for protein fitness optimization. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Additional Methods
A.1 Fitness predictor and oracle
The fitness predictors $g_{\phi}$ and $g_{\tilde{\phi}}$, as well as the oracle $g_{\psi}$ are directly re-used from (Kirjner et al., 2023) without any modifications. The authors use the identical 1D CNN architecture with 157k learnable parameters for all networks, which maps a one-hot encoded protein sequence to a scalar value representing the predicted fitness. Our method VLGPO allows to replace all of these components by re-trained models or alternative architectures, although it is known that simple CNNs can be competitive in such data-scarce settings (Dallago et al., 2021).
Since (Kirjner et al., 2023) does not explicitly report the final performance of the in-silico oracle, we compute the Mean Squared Error (MSE) on a subset of 512 randomly selected samples from the ground truth data set $\mathcal{S}^{*}$ to estimate its reliability. The oracle’s predictions closely follow the target fitness values, resulting in MSE values of 0.012240 for GFP and 0.002758 for AAV.
@@ A.2 Metrics
For the evaluation of the generated sequences, we use median fitness, diversity and novelty as described in  (Kirjner et al., 2023; Jain et al., 2022). The sequences are assessed by the oracle $g_{\psi}$ that was trained on the full data set $\mathcal{S}^{*}$ with minimum and maximum fitness values $y_{\mathrm{min}}$ and $y_{\mathrm{max}}$. The median normalized fitness of sampled sequences $x\in\mathcal{X}$ is computed using
@@ B.2 Variational Autoencoder
We train a VAE for each task to achieve at least $80\%$ reconstruction accuracy on a validation subset, balancing the latent prior with $\beta$. Table 7 shows these results. The lower accuracy in AAV compared to GFP is due to larger effects of mutations on shorter proteins (sequence length $d$). Since the data sets (see Table 2) are very small, the Kullback-Leibler (KL) divergence in Equation 4 is crucial to ensure a Gaussian latent distribution. We can also sample from the VAE and evaluate predicted fitness via $g_{\psi}$ (Table 7), although median normalized fitness is lower than in Table 2. This is expected, given the sparse latent space and the “hole” problem of VAEs (Rezende & Viola, 2018).
@@ B.4 Sampling Parameters
The determination of the hyperparameters $\alpha_{t}$ and $J$ in VLGPO sampling is performed through a grid search across different tasks. We keep $\alpha_{t}$ constant for all steps $t$. In addition to LABEL:fig:sampling, the resulting heatmaps for these two parameters on diversity are shown in LABEL:fig:sampling_sup. Overall, combining the findings on predicted fitness and diversity it seems to be a heuristic tradeoff to balance these metrics, unlike for GFP (medium), which allows to choose both parameters high. Likewise, for AAV (hard) and (medium) there appears to be a broad range from which suitable parameter choices for $\alpha_{t}$ and $J$ can be made. Only GFP (hard) displays substantially compromised diversity very quickly (note the limited numerical range displayed in the colorbar), hence we choose significantly lower values $\alpha_{t}=0.02$ and $J=5$. For the other tasks, we use the hyperparameter settings obtained from the grid search experiments, i.e., $\alpha_{t}\in\{0.97,1.2,0.56\}$ and $J\in\{39,19,37\}$ for AAV (medium), AAV (hard) and GFP (medium). Nevertheless, we do not observe major differences if these hyperparameters are adjusted slightly.
@@
Figure 13: Median fitness (left) and diversity (right) for all four tasks depending on employed ODE steps $K$ in sampling.